find an equation of the line passing through the given points
use function notation to write the equation
(-1,8) and (-2,13)
1
Expert's answer
2012-07-03T11:11:29-0400
(y-8)/(13-8)=(x-(-1))/(-2-(-1))
(y-8)/5=(x+1)/(-1)
8-y=5(x+1)
y=y(x)=-5x+3